OpenStudy (anonymous):

is the square root of 0.04 irrational or rational

OpenStudy (anonymous):

rational

OpenStudy (anonymous):

kool thanks

OpenStudy (sonic):

0.04 = 4/100 = 1/25 = \[\sqrt{1}/\sqrt{25}=1/5\] -> Rational.